OpenStudy (anonymous):

Mark deposits $700 each month in a retirement plan paying 6% compounded monthly. How much will he have in the account after 24 years

OpenStudy (john_es):

You should use the formula for compund interest, \[C=C_0\left(1+\frac{r}{100}\right)^n\] Where C_0 is the initial capital, r is the mensual rate in % and n is the time in months.

OpenStudy (tkhunny):

Build it!!!! i = 0.06 -- Annual interest rate to be compounded monthly. j = i/12 = 0.005 -- Monthly interest rate r = 1 + j = 1.005 -- Monthly Accumulation factor Now, just write ti down. 700(r + r^2 + r^3 + ... + r^(24*12)) = \(700\dfrac{r-r^{289}}{1-r}\)

OpenStudy (john_es):

Yes, a geometric progression is funnier to build.

OpenStudy (anonymous):

im getting 831.99 and its wrong

OpenStudy (tkhunny):

How did you get that? john_ES's formula is good for a single deposit. I built a complete formula. Please demonstrate your intermediate results. \(r^{289}\;=\;??\)

OpenStudy (anonymous):

Find the amount of the annuity if the deposit is $500 quarterly for 6 years at 7% compounded quarterly.

OpenStudy (tkhunny):

This is the EXACT same problem. There is no difference in the solution. Note: You really have to specify WHEN the payment is made. Beginning? End? Middle? Spread? What?

OpenStudy (anonymous):

That's the whole question.. there is nothing more to it....

OpenStudy (tkhunny):

I understand that, but it requires assumptions. Insist that there are better questions. They are often missing pieces. Very annoying.